discuss the refraction of light from a rarer medium to denser medium at a convex refracting surface
1
Expert's answer
2020-12-31T08:07:51-0500

As per the law of re-fraction of light, when ray of light passing from one medium to another medium then it will either move towards the normal or move away from the normal it depends on the medium.

It always follows the snell's law.

"\frac{\sin i}{\sin r}=\frac{\mu_2}{\mu_1}"

Here i and r the angle of incidence and angle of refraction and "\mu_1" and "\mu_2" are the refractive index of the medium.

We know that, the value of refractive index for the denser medium is greater than the value of refractive index of the rarer medium hence

"\mu_2>\mu_1"

So "i>r"

Hence, we can conclude that if ray of light passing from the rarer medium to denser medium then ray of light will bent towards the normal. As here we are using the convex refraction medium hence the distance of the image will get increase. and image will be magnified.